how do you solve n÷7+3=(n+2)÷7+-5
Answer by unlockmath(1688)

Hello,
With this: n÷7+3=(n+2)÷7+-5 let's get rid of the denominator by multiplying by 7 to result in: (Note: I'm reading this as (n÷7)+3=[(n+2)÷7]+-5
n+21=n+2-35 There's no solution
Now if it's written this way n÷(7+3)=(n+2)÷(7+-5) it will be:
n/10=(n+2)/2 Multiply both sides by 10 to get:
n=5n+10 Subtract 5n from both sides to get:
-4n=10 Divide both sides by -4 to get:
n=-10/4
There we go.
RJ
Check out a book I wrote at:
www.math-unlock.com